# Demystifying Behavior-Based Malware Detection at Endpoints

### v1

@@ 4.2. Machine Learning Details
We experiment with three ML approaches: (i) a bag-of-n-grams-based model (NGR), (ii) a hybrid model that combines convolution and attention (HYB), and (iii) a self-attention-based sequence model (ATT). These three approaches cover a wide range of designs proposed by prior work in behavior-based malware detection with ML, respectively, MalDy (karbab2019maldy, ), Neurlux (jindal2019neurlux, ) and, most recently, Nebula (trizna2023nebula, ). Please find the details on these models in Appendix B. We address the class imbalance in our training sets by oversampling the underrepresented classes (malware in EP and benign in SB datasets) (chawla2002smote, ). Note that our goal is not to develop a better model architecture or feature processing routine but to evaluate the effectiveness of state-of-the-art architectures at endpoint malware detection. We make our ML improvements without changing the fundamentals of these existing approaches.
@@
5. Endpoint Detection Evaluations
@@
Table 4 presents the endpoint detection performance of ML models (based on the methodology in Section 4.2), using three different data sets (SB1, SB2, and EP) for training and testing, following the scenarios in Table 2. Here, for example, SB1->SB2 and SB1->EP refer to the scenarios where the models are trained on the SB1 training set and evaluated on the SB2 and EP testing sets, respectively.
5.1. Evaluating the SB->EP Scenario
@@
First, we see significant gaps between SB->SB and SB->EP scenarios across the board. This gap is the largest when models are trained and ranked on the same sandbox, e.g., 95.0% vs. 11.2% for NGR trained and ranked on SB1. When models are ranked on a different sandbox, it often shrinks, e.g., 94.0% vs. 13.9% for NGR trained on SB1 and ranked on SB2. Nevertheless, the gap is still substantial even when we rank the models on their EP performance, e.g., 93.2% vs. 16.7% for NGR trained on SB1 and ranked on EP. Although there is a gap between the performances on different sandboxes, e.g., 94.0% vs. 70.4% for NGR trained on SB1 and ranked on SB2, it is significantly smaller than the SB->EP gap. Overall, this gap persists regardless of the training sandbox or the model type, hinting at the limitations of sandbox-trained models for endpoint detection. Notably, this gap is larger for models that perform better on the test traces from a sandbox that also produced the training set. In consequence, studies that rely on a single sandbox for evaluation in SB->SB (miller2016reviewer, ; karbab2019maldy, ; kuchler2021does, ; trizna2023nebula, ) are at a higher risk of producing worse models for SB->EP.
Table 5. Ranking correlations of sandbox-trained models according to their performances (TPR) on different test sets.
Next, we verify the possibility of finding a better model for SB->EP by ranking the models using traces from an unseen sandbox (not used for training). We rank our models (100+ in each setting) based on their SB1, SB2, and EP performances (TPR) and compute Spearman’s correlation coefficients between these rankings. Table 5 reveals that (except for one setting) (i) the rankings based on the training sandbox correlate poorly (sometimes negatively) with EP rankings, (ii) rankings based on an unseen sandbox correlate more strongly with EP rankings. For example, for ATT trained on SB1, the rankings based on SB1 and SB2 have $0.03$ and $0.45$ correlation with the rankings based on EP, respectively. We believe this is because a model that performs better on the traces from an unseen sandbox is less likely to have overfitted to features specific to the training sandbox. We discuss such features in more detail in Section 7. As a result, to get higher performance in SB->EP, we recommend performing model selection using traces from a different, unseen sandbox.
@@
Finally, we experiment with combining traces from two sandboxes (SB1 and SB2) for training. However, this brings only minor improvements for SB->EP over training on traces from a single sandbox. When we use EP traces to rank the models, training on both sandboxes achieves at most 17.5% TPR on the EP test set for NGR (11.7% for HYB, and 13.9% for ATT). In Section 7, we show that the diversity between the traces of a sample from two sandboxes is much lower than between its traces from two endpoint hosts. This lack of diversity makes combining SB1 and SB2 traces ineffective for increasing the EP performance.
@@
Applying Drift Detection. We study the low performance in SB->EP through the lens of distribution shift, which occurs if the process generating inputs, i.e., program traces, changes between the training and testing phases. Concept drift is a highly-studied type of distribution shift in malware detection that stems from malware evolution (pendlebury2019tesseract, ). In past work on malware concept drift, the distribution shift builds up over time in concept drift, e.g., over 50% performance drop in 2 years (barbero2022transcending, ). In contrast, we find that in SB->EP the shift manifests immediately after the model is deployed.
@@
One way to combat malware concept drift is to use a drift detector (DD) to identify samples that might have drifted and abstain from classifying them (jordaney2017transcend, ; barbero2022transcending, ; yang2021cade, ; chen2023continuous, ). DDs aim to prevent mistakes by rejecting out-of-distribution (OOD) samples that diverge from the training distribution while accepting in-distribution (ID) samples on which predictions are reliable (jordaney2017transcend, ; yang2021cade, ). We apply drift detection to quantify the distribution shift present in SB->EP. We employ a modern technique ReAct (sun2021react, ) to assign an outlierness score to each sample, among which the highest-scoring ones are rejected. ReAct rectifies a model’s penultimate layer activations, which is shown to make the prediction probabilities of a model a more calibrated indicator of outlierness. We selected ReAact due to its flexible approach that is agnostic to the underlying model as long as it’s a neural network.
Figure 2. The rejection rates resulting from drift detection.
@@
Finally, in Table 6, we present the TPR on the accepted (remaining) traces in EP and SB2 for increasing values of $K$. Rejecting the traces in EP or SB2 increases the performance, but this increase is less pronounced for EP, e.g., gaining $\sim$10TPR (over the baselines in Table 4) takes rejecting 67% and 22% of malware traces in EP and SB2, respectively. This also distinguishes the distribution shift in SB->EP from concept drift; rejection rates must be high for meaningful performance gains due to the magnitude of the shift.
($K$) ID RejSB1->EP SB1->SB2 TPR MalBenTPR MalBen5% 20.2% 31.1% 20.5% 79.2% 22.0% 21.5% 10% 23.5% 53.2% 39.0% 89.4% 38.2% 36.5% 15% 25.6% 67.7% 46.4% 88.3% 52.4% 53.6% 20% 42.6% 72.2% 49.0% 87.7% 60.8% 60.6% 25% 53.1% 78.5% 52.5% 82.9% 68.5% 70.8%
@@
Table 6. The TPR of an NGR model in SB1->SB2 and SB1->EP scenarios, with increasing ID rejection rates, and the corresponding rejection rates of malware and benign traces.
5.2. Evaluating the EP->EP Scenario
@@
For our models in EP->EP, we select up to 4 EP traces per sample to form our training set. As we will show in Section 7, the behavior variability among the real-world traces of a sample (avllazagaj2021malware, ) impacts the model’s predictions. Consequently, including more traces per sample for training provides better coverage of the behaviors in the wild. We find that using more than 4 traces per sample provides diminishing returns, e.g., for 1, 2, 4, and 16 traces per sample, the TPR of NGR is 43.0%, 44.8%, 49.5%, and 49.9%, respectively.
@@
In Table 4, we see a much higher EP performance in EP->EP than in SB->EP, as expected. Although training and testing sets are from similar distributions in EP->EP (as opposed to SB->EP), the performance still lags behind the SB->SB performance, e.g., 95% vs. 50% TPR and 99% vs. 88% AUC (Table 15). In the following sections, we shed light on this discrepancy and what makes endpoint detection particularly more challenging than a sandbox-only scenario.
6. Endpoint Detection Deals With Difficult-To-Classify Samples
@@
As described in Section 3.1, an endpoint detector operates on samples that cannot be classified statically—a subset of all samples that exist in the wild. The filtering effect this introduces on the distribution of samples relevant to endpoint detection has not been systematically measured before. To illustrate, we explore (based on Table 13) the top malware families in our EP test set (undetected samples) and the SB test set (samples from public corpora) Although both datasets include only the samples first seen in 2017-2018, they encode vastly different priors over samples. For example, generic malware (i.e., malware that could not be placed into any known family) covers 23.8% of the EP samples vs. 4.8% of the SB samples. In the EP set, there are families such as Khalesi and Emotet that were rampant in 2018 (emotet, ). Conversely, in the SB set, we see older families such as Sivis and Upatre (circa 2013-2014), whose variants still spread to this day. Moreover, in Table 13, we share top benign sample publishers (extracted from the code-signing certificates when available), indicating a similar distribution difference in benign samples. Overall, the typical distribution of training and testing samples used by prior work in SB->SB (karbab2019maldy, ; jindal2019neurlux, ; dambra2023decoding, ) is substantially different than the distribution of testing samples in SB->EP and EP->EP.
@@
Borderline Samples and Label Noise. A common practice in related work in malware detection with ML is discarding borderline samples on which the anti-malware engines deployed by VirusTotal are not in consensus. Different works use different criteria; for example, they discard a sample if it is detected by less than 20 engines (mantovani2020prevalence, ) or five engines (kuchler2021does, ), or by more than 0 and less than 40 engines (lucasadversarial, ). Even for works that do not explicitly discard borderline samples, their data collection results in few of them. For example, recent related work has suggested that dynamic features have limited benefits over static features (dambra2023decoding, ). The benign samples used in this work were from a trusted Windows software repository, and 93% of their malware samples were detected by 20+ engines.
@@
Based on prior definitions, we consider samples detected by more than 0 and less than 20 engines borderline. This results in 27.7% borderline samples in our EP test set (vs. 15.9% in the SB test set). The prior practice would have discarded over a quarter of the samples that a real-world endpoint detector encountered. Next, we will show that this artificially inflates the measured success of a model. As described in Section 4.1, we do not discard any samples and label a sample as malware if more than five engines detect it. This threshold poses a trade-off for the ground truth: as it increases, more malware samples may be labeled as benign, and as it decreases, more benign samples may be labeled as malware. We present a case study in Section 6.2 to demonstrate how this trade-off plays out.
6.1. The Impact of Sample Distributions
@@
Here, we measure how the distribution of samples in the test set changes a model’s measured performance. To this end, we resample our test sets according to some criteria, e.g., following the malware family distribution in the EP test set, and evaluate our models on these new test sets. This simulates different distributions over a model’s test samples and allows us to do controlled experiments. We focus on NGR models due to their superior EP performance in Section 5.
EX# SB1->SB1 SB1->EP EP->EP
TPR AUC TPR AUC TPR AUC
EX0 No Resampling - Orig. Distributions
93.2 99.0 16.7 78.4 49.5 87.5
EX1 Orig. Distributions w/o Generic Malware
95.6 99.5 18.6 81.4 56.8 91.4
EX2 Only Generic Malware
38.4 85.8 9.8 67.8 23.9 74.2
EX3 Discard Borderline Samples
96.7 99.5 22.6 83.9 66.9 94.2
EX4 Malw. Resampled Following EP Test (w/ Generic)
63.2 94.2 16.7 78.4 49.5 87.5
EX5 Malw. Resampled Following EP Test (w/o Generic)
72.1 97.2 18.6 81.4 56.8 91.4
@@
Table 7. The impact of sample distributions in the test set on model evaluations. Each experiment is denoted as EX0-5.
@@
Table 7 presents the results of our resampling experiments. We resample the test sets according to five criteria—denoted as EX0--5 where EX0 is testing on the original test set of each respective scenario. In EX1, leaving out generic malware samples from original test sets leads to a significant boost in performance, whereas keeping only generic malware in EX2 causes a massive drop. This aligns with prior claims that generic malware is at the borderline and much harder to classify (ugarte2019close, ). The fact that Generic has 5$\sim$ 5$\times$ 5 × higher coverage in the EP test set than in the SB test set (23.8% vs. 4.8%) also demonstrates the filtering effect that funnels more difficult samples to the endpoint detector. Next, in EX3, we discard the borderline samples (identified in the previous section) from all test sets, simulating a common prior practice (kuchler2021does, ; mantovani2020prevalence, ; dambra2023decoding, ). This yields a significant boost in all scenarios: 5.9% TPR boost in SB1->EP and 17.4% in EP->EP. In EX4 and EX5, we resample the test sets to match the distribution of malware families in the EP test set, with and without generic malware, respectively. Note that, for SB1->EP and EP->EP, EX4 is equivalent to EX0 and EX5 is equivalent to EX1. We observe a great decrease in the original SB1->SB1 performance, e.g., from 93.2% TPR to 63.2% in EX4 and to 72.1% in EX5. These results suggest that the difference in distribution of malware families explains some of the performance drop when applying these classifiers to EP traces, perhaps because EP traces contain more harder-to-classify malware families, but there remains a large gap that is not explained by differences in the distribution of families. Despite neither including any generic malware, the SB1->SB1 performance in EX5 is significantly lower than EX1. This suggests that non-generic malware samples relevant to endpoint detection are generally more difficult to classify correctly, whether based on their sandbox or endpoint traces. Overall, these experiments reveal how the performance of prior behavior-based detectors may be over-estimated because they are evaluated on samples that are not representative of the distribution of samples faced in endpoint detection.
6.2. Case Studies

### v2

@@ IV-B Machine Learning Details
We experiment with three deep-learning-based approaches: (i) a bag-of-n-grams-based model (NGR), (ii) a hybrid model that combines convolution and attention (HYB), and (iii) a self-attention-based sequence model (ATT). These approaches cover a wide range of designs proposed by prior work in behavioral malware detection with ML, respectively, MalDy [5], Neurlux [6] and, most recently, Nebula [7] (see Appendix B for details on these models). We address the class imbalance in our training sets by oversampling the underrepresented classes (malware in EP and benign in SB datasets). Note that our goal is not to develop a new architecture or feature processing routine but to evaluate the state-of-the-art ones at endpoint malware detection. We make our ML improvements without changing the fundamentals of existing approaches.
@@
V Endpoint Detection Evaluations
@@
Table IV presents the detection performance of ML models (based on the methodology in Section IV-B), using three different data sets (SB1, SB2, and EP) for training and testing, following the scenarios in Table II. Here, for example, SB1->SB2 and SB1->EP refer to the scenarios where the models are trained on the SB1 training set and evaluated on the SB2 and EP testing sets, respectively. See Table XIII for the AUC results.
V-A Evaluating the SB->EP Scenario
@@
We see wide performance gaps between SB->SB and SB->EP scenarios. The gap widens when models are trained and selected using the same sandbox, e.g., 95.0% vs. 11.2% for NGR trained and selected using SB1. When models are selected using a different sandbox, the gap shrinks, e.g., 94.0% vs. 13.9% for NGR trained on SB1 and selected using SB2. The gap is still substantial even when we select the models based on their EP performance: e.g., 93.2% vs. 16.7% for NGR trained on SB1 and selected using EP. Although there is a gap between a model’s performances on different sandboxes, e.g., SB1->SB1 vs. SB1->SB2, it is smaller than the SB->EP gap. Overall, this gap persists regardless of the training sandbox or the model type, hinting at the limitations of sandbox-trained models for endpoint detection.
We observed that the gap widens when model training and selection use traces from the same sandbox. In Appendix C, we perform an experiment showing this is a generally valid observation. We believe this is because a model that performs better on the traces from an unseen sandbox is less likely to have overfitted to features specific to the training sandbox. We discuss such features in more detail in Section VII. This implies that studies that rely on a single sandbox for training and evaluation in SB->SB [55, 5, 28, 7] are at a higher risk of producing worse models for SB->EP. Consequently, we recommend performing model selection in SB->EP using traces from a different, unseen sandbox.
@@
Finally, we explore combining traces from two sandboxes (SB1 and SB2) for training. However, this brings only minor improvements for SB->EP over training on traces from one sandbox. When we use EP traces for model selection, training on both sandboxes achieves at most 17.5% TPR on the EP test set for NGR (11.7% for HYB, and 13.9% for ATT). In Section VII, we show that the diversity between the traces of a sample from two sandboxes is much lower than between its traces from two endpoint hosts. This lack of diversity makes combining SB1 and SB2 traces ineffective for increasing the EP performance.
The Impact of Trace Standardization. The standardized trace format (see Section IV-A) removes features from sandbox traces unavailable in endpoint traces. In Section A-B, to understand whether this might hurt the SB->EP performance, we experiment with training and testing models on unstandardized sandbox traces. This shows that standardization moderately hurts the in-domain performance (test traces from the training sandbox) while improving generalization to other domains (traces from other sandboxes). Consequently, we believe standardization helps with improving out-of-domain performance in SB->EP.
@@
Applying Drift Detection. We study the low performance in SB->EP through the lens of distribution shift, which occurs if the process generating inputs, i.e., program traces, changes between the training and testing phases. Concept drift is a highly-studied type of distribution shift in malware detection that stems from malware evolution [38]. In past work on malware concept drift, the distribution shift builds up over time in concept drift, e.g., over 50% performance drop in 2 years [25]. In contrast, we find that in SB->EP, the shift manifests immediately after the model is deployed.
@@
One way to combat malware concept drift is to use a drift detector (DD) to identify samples that might have drifted from the training distribution [24, 25, 60, 39]. The classifier then rejects such out-of-distribution (OOD) samples while accepting the in-distribution (ID) ones on which predictions are reliable [24, 60]. We apply drift detection to quantify the distribution shift present in SB->EP. We employ a recent technique ReAct [61] to assign an outlierness score to each sample, among which the highest-scoring ones are rejected. ReAct rectifies a model’s penultimate layer activations, which calibrates the prediction probabilities of a model to be a better indicator of outlierness. We selected ReAact due to its flexible approach, which can be applied to any neural network model.
We experiment with a NGR model trained on the SB1 training set. We tune ReAct to reject $K\in[0,15]$of the traces in the SB1 test set (i.e., the ID rejection rate) and separately measure for EP and SB2 test sets (i) corresponding rejection rates on the malware and benign traces; and (ii) the TPR on the accepted traces. Table V presents the results.
@@
Regarding (ii), classification with rejection increases the performance for both EP and SB2 test sets. However, this increase is less pronounced for EP, e.g., gaining $\sim$10TPR (over the baselines in Table IV) takes rejecting 67% and 22% of malware traces in EP and SB2, respectively. This also distinguishes the distribution shift in SB->EP from concept drift; rejection rates must be high for meaningful performance gains due to the magnitude of the shift.
($K$) ID RejSB1->EP SB1->SB2 TPR MalBenTPR MalBen5% 20.2% 31.1% 20.5% 79.2% 22.0% 21.5% 10% 23.5% 53.2% 39.0% 89.4% 38.2% 36.5% 15% 25.6% 67.7% 46.4% 88.3% 52.4% 53.6%
@@
TABLE V: The TPR of an NGR model in SB1->SB2 and SB1->EP scenarios, with increasing ID rejection rates and the corresponding rejection rates of malw. and benign traces.
V-B Evaluating the EP->EP Scenario
@@
For our models in EP->EP, we select up to 4 EP traces per sample to form our training set. As we will show in Section VII, the behavior variability among the real-world traces of a sample [15] impacts the model’s predictions. Consequently, including more traces per sample for training provides better coverage of the behaviors in the wild. We find that using more than 4 traces per sample provides diminishing returns, e.g., for 1, 2, 4, and 16 traces per sample, the TPR of NGR is 43.0%, 44.8%, 49.5%, and 49.9%, respectively.
@@
In Table IV, we see a much higher EP performance in EP->EP than in SB->EP, as expected. Although training and testing sets are from similar distributions in EP->EP (as opposed to SB->EP), the performance still lags behind the SB->SB performance, e.g., 95% vs. 50% TPR and 99% vs. 88% AUC (Table XIII). In the following sections, we shed light on this discrepancy and what makes endpoint detection particularly more challenging than a sandbox-only scenario.
Takeaways: Distribution shift causes ML methods trained on sandbox traces to perform poorly on endpoint traces (70+TPR drop). Even when trained on endpoint traces, these methods’ endpoint performance still cannot reach their sandbox-based performance ($\sim$20vs. 95% TPR), hinting at the inherent challenges in endpoint detection.
VI Endpoint Detection Deals With Difficult-To-Classify Samples
@@
As described in Section III-A, an endpoint detector operates on samples that cannot be classified statically—a subset of all samples in the wild. The filtering effect this introduces on the distribution of samples relevant to endpoint detection has not been systematically measured before. To illustrate, we examine (based on Table XIV) the top malware families in our EP test set (undetected samples) and the SB test set (the samples from public corpora). Although both datasets include only the samples first seen in 2017-2018, they encode different priors over samples. For example, generic malware (i.e., malware that could not be placed into any known family) covers 23.8% of the EP samples vs. 4.8% of the SB samples. In the EP set, there are families such as Khalesi and Emotet that were rampant in 2018 [62]. Conversely, in the SB set, we see older families such as Sivis and Upatre (circa 2013-2014), whose variants still spread to this day. Moreover, in Table XIV, we share top benign sample publishers (extracted from the code-signing certificates when available), indicating a similar distribution difference in benign samples. Overall, the typical distribution of training and testing samples used by prior work in SB->SB [5, 6, 26] is substantially different than the distribution of testing samples in SB->EP and EP->EP.
@@
Borderline Samples and Label Noise. A common practice in related work in malware detection with ML is discarding borderline samples on which the anti-malware engines on VirusTotal are not in consensus. Different works use different criteria; for example, they discard a sample if it is detected by more than zero and less than five engines [28], 20 engines [63] or 40 engines [64]. Even works that do not explicitly discard borderline samples may follow a biased data collection process. For example, recent related work has suggested that dynamic features have limited benefits over static features [26]. The benign samples used in this work were from a trusted Windows software repository, and 93% of their malware samples were detected by 20+ engines.
@@
We consider the samples detected by more than zero and less than 20 engines borderline. This results in 27.7% borderline samples in our EP test set (vs. 15.9% in the SB test set). The prior practice would have discarded over a quarter of the samples that a real-world endpoint detector encountered. Next, we will show that this artificially inflates the measured success of a model. As described in Section IV-A, we do not discard any samples and label a sample as malware if more than five engines detect it. This threshold poses a trade-off for the ground truth: as it increases, more malware samples may be labeled as benign, and as it decreases, more benign samples may be labeled as malware. We present a case study in Section VI-B to demonstrate how this trade-off plays out.
VI-A The Impact of Sample Distributions
@@
Here, we measure how the distribution of samples in the test set changes a model’s measured (and perceived) performance. To this end, we resample our test sets according to some criteria, e.g., following the malware family distribution in the EP test set, and evaluate our models on these new test sets. This simulates different distributions over a model’s test samples and allows us to do controlled experiments. We focus on NGR models due to their superior EP performance in Section V.
EX# SB1->SB1 SB1->EP EP->EP TPR AUC TPR AUC TPR AUC EX0 No Resampling - Orig. Distributions 93.2 99.0 16.7 78.4 49.5 87.5 EX1 Orig. Distributions w/o Generic Malware 95.6 99.5 18.6 81.4 56.8 91.4 EX2 Only Generic Malware 38.4 85.8 9.8 67.8 23.9 74.2 EX3 Discard Borderline Samples 96.7 99.5 22.6 83.9 66.9 94.2 EX4 Malw. Resampled Following EP Test (w/ Generic) 63.2 94.2 16.7 78.4 49.5 87.5 EX5 Malw. Resampled Following EP Test (w/o Generic) 72.1 97.2 18.6 81.4 56.8 91.4
@@
TABLE VI: The impact of sample distributions in the test set on model evaluations. EX0-5 denote different experiments.
@@
Table VI presents the results of our resampling experiments. We resample the test sets according to five criteria—denoted as EX0--5 where EX0 is testing on the original test set of each respective scenario. In EX1, leaving out generic malware samples from original test sets leads to a significant boost in performance, whereas keeping only generic malware in EX2 causes a massive drop. This aligns with prior claims that generic malware is at the borderline and harder to classify [54]. The fact that Generic has 5$\sim$ 5$\times$ 5 × higher coverage in the EP test set than in the SB test set (23.8% vs. 4.8%) also demonstrates the filtering effect that funnels more difficult samples to the endpoint detector. Next, in EX3, we discard the borderline samples (identified in the previous section) from all test sets, simulating a common prior practice [28, 63, 26]. This yields a significant boost in all scenarios: 5.9% TPR boost in SB1->EP and 17.4% in EP->EP. In EX4 and EX5, we resample the test sets to match the distribution of malware families in the EP test set, with and without generic malware, respectively. Note that, for SB1->EP and EP->EP, EX4 is equivalent to EX0 and EX5 is equivalent to EX1. We observe a large decrease in the original SB1->SB1 performance, e.g., from 93.2% TPR to 63.2% in EX4 and to 72.1% in EX5. Despite neither including any generic malware, the SB1->SB1 performance in EX5 is significantly lower than EX1. This suggests that non-generic malware samples relevant to endpoint detection are generally harder to classify correctly, whether based on their sandbox or endpoint traces. Consequently, the performance of prior behavioral detectors may be overestimated as they are evaluated on samples that are not representative of the distribution of samples faced in endpoint detection. Although the distribution of malware samples explains most of the performance loss when ML classifiers are applied to EP traces, an unexplained gap remains (63.2% vs. 49.5%), which we attempt to explain in Section VII.
VI-B Case Studies